# Automatic Speech Recognition using Advanced Deep Learning Approaches: A survey

### v1

@@ Abstract
Recent advancements in deep learning (DL) have posed a significant challenge for automatic speech recognition (ASR). ASR relies on extensive training datasets, including confidential ones, and demands substantial computational and storage resources. Enabling adaptive systems improves ASR performance in dynamic environments. DL techniques assume training and testing data originate from the same domain, which is not always true. Advanced DL techniques like deep transfer learning (DTL), federated learning (FL), and reinforcement learning (RL) address these issues. DTL allows high-performance models using small yet related datasets, FL enables training on confidential data without dataset possession, and RL optimizes decision-making in dynamic environments, reducing computation costs.
This survey offers a comprehensive review of DTL, FL, and RL-based ASR frameworks, aiming to provide insights into the latest developments and aid researchers and professionals in understanding the current challenges. Additionally, transformers, which are advanced DL techniques heavily used in proposed ASR frameworks, are considered in this survey for their ability to capture extensive dependencies in the input ASR sequence. The paper starts by presenting the background of DTL, FL, RL, and Transformers and then adopts a well-designed taxonomy to outline the state-of-the-art (SOTA) approaches. Subsequently, a critical analysis is conducted to identify the strengths and weaknesses of each framework. Additionally, a comparative study is presented to highlight the existing challenges, paving the way for future research opportunities.
@@ 1.1 Preliminary
Advancements in artificial intelligence (AI) have significantly improved human-machine interaction (HMI), especially with technologies that convert speech into executable actions. automatic speech recognition (ASR) emerges as a leading communication technology in HMI, extensively utilized by corporations and service providers for facilitating interactions through AI platforms like chatbots and digital assistants. Spoken language forms the core of these interactions, emphasizing the necessity for sophisticated speech processing in AI systems tailored for ASR. ASR technology encompasses the analysis of (i) acoustic, lexical, and syntactic aspects; and (ii) semantic understanding. The acoustic model (AM) processing includes speech coding [1], enhancement [2], and source separation [3], alongside securing speech via steganography [4, 5, 6] and watermarking [7, 8, 9]. These components are integral to audio analysis. On the other hand, the semantic model (SM), often identified as language model (LM) processing in literature, involves all natural language processing (NLP) techniques. This AI branch aims at teaching computers to understand and interpret human language, serving as the basis for applications like music information retrieval [10], sound file organization [11], audio tagging, and event detection (ED) [12], as well as converting speech to text and vice versa [13], detecting hate speech [14], and cyberbullying [15]. Employing NLP across various domains enables AI models to effectively comprehend and respond to human inputs, unveiling extensive research prospects in diverse sectors.
@@
S2S
sequence-to-sequence
@@
GAN
generative adversarial network
MTL
multitask learning
FMTL
federated multi-task learning
DSLM
domain-specific language modeling
LLM
large language model
DDQN
double deep Q-network
AC
actor-critic
SARSA
State–action–reward–state–action
DDPG
deep deterministic policy gradien
Recent advancements in ASR have been significantly propelled by the evolution of deep learning (DL) methodologies. An extensive range of DL models has been developed, demonstrating remarkable improvements and surpassing former SOTA achievements [16, 17]. Transformers, a notable innovation within these DL approaches, have become a cornerstone in advancing various NLP tasks, including ASR. Initially conceptualized for sequence-to-sequence (S2S) applications in NLP, their success is largely attributed to their adeptness at discerning long-range dependencies and complex patterns within sequential data. A hallmark of transformer models is their utilization of an attention mechanism, which precisely focuses on specific portions of the input sequence during prediction tasks. This mechanism is particularly effective in ASR, facilitating the detailed modeling of contextual nuances and the interconnections among acoustic signals, essential for accurate transcription. Models such as the Transformer Transducer, Conformer, and ESPnet, leveraging self-attention and parallel processing, have achieved leading performance in ASR tasks. Their robustness across diverse languages further underscores their capability to adapt to a wide range of linguistic features and acoustic variations, making transformers an exceptionally promising option for enhancing ASR systems, surpassing the constraints of conventional models.
@@ 1.2 Contribution of the paper
In recent years, numerous survey papers have been published to assess various aspects of ASR models. Some of these surveys concentrate on specific languages, such as Portuguese [23], Indian [24], Turkish [25], Arabic [26] and tonal languages (including Asian, Indo-European and African) [27]. Additionally, Abushariah et al.’s review emphasizes bilingual ASR [28]. On the non-specific language review front, specific areas within ASR have been targeted, including ASR using limited vocabulary [29], ASR for children [30], error detection and correction [31], and unsupervised ASR [32]. Systematic reviews with a focus on neural networks [33] and deep neural networks [34] have also been proposed. In another comprehensive review, Malik et al. [35] discussed diverse feature extraction methods, SOTA classification models, and some deep learning approaches. Recently, the authors presented an ASR review focused on DTL for ASR [36]. Table 1 presents a summary of the main contributions of the proposed ASR review compared to other existing ASR reviews/surveys.
This survey article offers several significant enhancements and additions compared to previous ASR surveys. Firstly, it consolidates works that utilize both ASR and advanced DL approaches, providing a comprehensive overview of their intersection. Secondly, it provides performance evaluation results of all considered approaches. Thirdly, it includes metrics and dataset reviews used in ASR models. Furthermore, it tackles ongoing challenges and consequently proposes future directions. The main contributions of this article can be summarized as follows:
@@
Table 1: Contribution comparison of the proposed contribution against other hand ASR review. The tick mark (✓) indicates that the specific field has been addressed, whereas the cross mark () means addressing the specific fields has been missed.
Refs Year
Description of the survey/review
Advanced DL methods
Performances
Metrics
Dataset
Current
Future
DRL
FL
TL
Transf.
evaluation
review
challe/Gaps
directions
[31] 2018
ASR review for error errors detection and correction
[34] 2019
Systematic review on DL-based speech recognition
[24] 2020
ASR survey for Indian languages
[23] 2020
ASR survey for Portuguese language
[25] 2020
ASR survey for Turkich language
[35] 2021
ASR survey
[27] 2021
ASR survey for tonal languages
[32] 2022
Unsupervised ASR review
[30] 2022
ASR Systematic review for children
[29] 2022
ASR survey for limited vocabulary
[26] 2022
ASR Systematic review for Arabic language
[28] 2022
Bilingual ASR review
[33] 2023
ASR survey on neural network techniques
[36] 2023
ASR based on DTL review
Our 2024
ASR review on advanced DL techniques
@@ 1.4 Structure of the paper
This paper is structured into six sections. The current section provides an introduction to the paper. Section 1.3 providing background on AM and LM, and reviewing evaluation metrics and datasets utilized in ASR. Moving forward, Section 4 delves into a comprehensive review of recent advancements in ASR utilizing advanced DL approaches, including Transformers, DTL, FL and DRL. Sections 5 and 6 respectively address the existing challenges and future directions concerning advanced DL-based ASR. Finally, Section 7 presents concluding remarks summarizing the key findings of the paper.
@@
table[t!] Literature acquisition databases. to be modified Database Research Articles Conference Papers Book Chapter Total ACM 5 1 6 Elsevier 26 26 Springer 23 10 33 IEEE 39 28 67 Others 44 29 70
2 Overview of DTL techniques for speech recognition
2.1 Taxonomy of existing DTL techniques
To date, there is no standardized and comprehensive technique for classifying DTL into categories. However, DTL algorithms could be classified into several types depending on what, when, and how knowledge is transferred:
(a) What knowledge is transferred? Enquires about which characteristics of knowledge are transferable across domains or tasks. Some information is particular to certain domains or tasks, while other knowledge is shared across domains and can aid increase performance in the target task or domain. Based on this definition, DTL could be feature-based, instance-based, relation-based, or model-based.
(b) How is knowledge transferred? Enquires about which learning algorithms must be implemented to transfer the knowledge.
(c) When is knowledge transferred? Inquires as to when and under what circumstances knowledge should or should not be transferred.
Furthermore, researchers have proposed taxonomies to categorize DTL-based ASR techniques. For instance, Niu et al. [21] present a taxonomy with two levels. The first level consists of four sub-groups based on the availability of labeled data and the data modality in the source and target domains. These sub-groups include inductive DTL, transductive DTL, cross-modality DTL, and unsupervised DTL [80]. Table 2 provides a summary of these possibilities. Moreover, each sub-group at the first level can be further divided into four distinct learning types: learning on instances, learning on features, learning on parameters, and learning on relations.
[…]
Domains Tasks
Math. propriety
Sub-categories / Usage
Traditional ML/DL $\mathbb{D}_{S}=\mathbb{D}_{T}$ $\mathbb{T}_{S}=\mathbb{T}_{T}$
$X_{S}\neq X_{T}$,
$Y_{S}=Y_{T}$
ASR model trained with $X_{S}$ database and used to recognise $X_{T}$ database.
Inductive DTL $\mathbb{D}_{S}\cong\mathbb{D}_{T}$ $\mathbb{T}_{S}\neq\mathbb{T}_{T}$
$X_{S}\neq X_{T}$,
$Y_{S}\exists,Y_{T}\exists$
If $Y_{S}\exists$, DTL is multitask learning. If $Y_{S}\nexists$, DTL is self-taught learning, thus $\chi_{S}\cong\chi_{T}$.
Transductive DTL $\mathbb{D}_{S}\neq\mathbb{D}_{T}$ $\mathbb{T}_{S}=\mathbb{T}_{T}$
$P(X_{S})\neq P(X_{T})$,
$Y_{S}\exists,Y_{T}\nexists$ ,
$\chi_{S}=\chi_{T}$
When $\chi_{S}=\chi_{T}$, DTL is is related to DA. If $\mathbb{D}_{T}\exists!$ and $\mathbb{T}_{T}\exists!$, DTL is used for sample selection bias or covariate shift.
Cross-modality DTL $\mathbb{D}_{S}\neq\mathbb{D}_{T}$ $\mathbb{T}_{S}\neq\mathbb{T}_{T}$
$P(Y_{S}/X_{S})\neq P(Y_{T}/X_{T})$,
$Y_{S}\neq Y_{T}$, $\chi_{S}\neq\chi_{T}$
i,e. the dataset $X_{S}$ of $\mathbb{D}_{S}$ is speech data, and the dataset $X_{T}$ of $\mathbb{D}_{T}$ is text data.
Unsupervised DTL $\mathbb{D}_{S}\varsubsetneq\mathbb{D}_{T}$ $\mathbb{T}_{S}\varsubsetneq\mathbb{T}_{T}$
$Y_{S}\nexists,Y_{T}\nexists$
DTL used for clustering, dimensionality reduction, and density estimation, etc.
Inductive DTL In comparison to classical ML, which may be used as a reference for DTL comparison, and given that the target tasks $\mathbb{T}_{T}$ are distinct from the source tasks $\mathbb{T}_{S}$, the goal of inductive DTL is to enhance the target prediction function $\mathbb{F}_{T}$ in the TD, mentioned above in subsection LABEL:sub12. However, the SD $\mathbb{D}_{S}$ and TD $\mathbb{D}_{T}$ may not always be the same (Table 2). The inductive DTL can be stated similarly to the following two cases, depending on whether labeled or unlabeled data is available:
(a) Multi-task DTL: The SD has a huge labeled database ($X_{S}$ labeled with $Y_{S}$), which is a distinctive form of multi-task learning. However, with multi-task approaches, many tasks $(T_{1},T_{2},\dots,T_{n})$ are learned at the same time (in parallel), including both source and target activities (tasks).
(b) Sequential DTL: ( Commonly known as self-taught learning) Dataset is not labeled in the SD ($X_{S}$ is not labeled with $Y_{S}$) but the labels are available in the destination domain ($X_{T}$ is labeled with $Y_{T}$). Sequential learning is a DL system that can be realized in two steps for classification purposes. The first step is the feature representation transfer, which is learned from a large collection of the unlabeled datasets, and the second stage is when this learned representation is applied to labeled data to accomplish classification tasks. Hence, sequential DTL is a method of sequentially learning a number of activities (Tasks). The spaces between the source and destination domains may differ. For example, let suppose we have a pre-trained model $M$ and consider applying DTL to a number of tasks $(T_{1},T_{2},\dots,T_{n})$. We learn a specific task $\mathbb{T}_{T}$ at each time step $t$, which is slower than multi-task learning. However, when not all the tasks are present during training time, it might be beneficial. Sequential DTL is additionally classified into several types [alyafeai2020survey]:
  • 1-
    Fine-tuning: The principle is to learn a new function $\mathbb{F}_{T}$ that translates the parameters $\mathbb{F}_{T}(W_{S})=W_{T}$ by using $M$, given a pre-trained model $M_{S}$ having $W_{S}$ as weights and target task $\mathbb{T}_{T}$ having $W_{T}$ as weights. The settings can be adjusted across all layers or just some of (Fig. 3 (a)). The learning rate for each layer could be distinct (discriminative fine tuning). A new set of parameters $K$ could be added to most of the tasks so that $\mathbb{F}_{T}(W_{T},K)=W_{S}\circ K$.
  • 2-
    Adapter modules: Given an $M_{S}$ model that has been pre-trained and output $W_{S}$, for a target task $\mathbb{T}_{T}$. The adapter module aims to lunch a different set of parameters $K$ that is too much less than $W_{S}$, i.e, $K\ll W_{S}$. $K$ and $W_{S}$ must have the ability to be decomposed into more compact modules such that, $W_{S}=\{w\}_{n}$ and $K=\{k\}_{n}$. The adapter module permit learning the following new function $\mathbb{F}_{T}$:
    $\mathbb{F}_{T}(K,W_{S})=k_{1}^{\prime}\circ w_{1}\circ\dots k_{n}^{\prime}\circ w_{n}.$ (1)
    According to equation 1, during the adaptation procedure, the set of original weights $W_{S}=\{w\}_{n}$ is left unaltered, but the set of weights K is changed to $K^{\prime}=\{k^{\prime}\}_{n}$. The principle of the adaptation domain is illustrated in Fig. 3 (b).
  • 3-
    Feature based: Interested only in learning concepts and representations, at various levels, such as, word, character, phrase, or paragraph embedding $E$. The collection of $E$ based on a model $M$ remains unaltered, i.e., $\mathbb{F}_{T}(W_{S},E)=E\circ W^{\prime}$, in the way that $W^{\prime}$ is fine-tuned. For example, researchers have applied the generative adversarial network (GAN) principle to DTL where, the generators send features from the SD and the TD to a discriminator, which determines the source of the features and feeds the result back to the generators until they can no longer be distinguished. In this procedure, GAN obtains the common properties of two domains, as shown in Fig. 3 (c).
  • 4-
    Zero-shot: Is the easiest method among all of the others. Making the assumption that the parameters $W_{S}$ can’t be modified or add $K$ as a new parameters to a pre-trained model $M_{S}$ using $W_{S}$. To put this into context, in zero-shot there is no training technique to optimize or learn new parameters.
Figure 3: Structures of: (a) Fine-tuning, (b) DA, and (c) DTL-based GAN.
2.1.1 Transductive DTL
Compared to the traditional ML, which can be considered as a reference for DTL comparison, and given that the TDs $\mathbb{D}_{T}$ are distinct from the SDs $\mathbb{D}_{S}$. The SD has a labeled dataset ($X_{S}$ labeled with $Y_{S}$), whereas the TD has no labeled dataset, the source and target tasks are equal ( Table 2). The goal of transductive DTL is to build the target prediction function $\mathbb{F}_{T}$ in the $\mathbb{D}_{T}$ by knowledge of the $\mathbb{D}_{S}$ and $\mathbb{T}_{T}$. Furthermore, the transductive DTL environment may be further classified into two categories depending on different conditions between the source and destination domains:
(a) Domain adaptation (DA): The feature spaces across domains, $\chi_{S}$ and $\chi_{T}$, are the identical, but the marginal probability distributions of the input dataset are not, $P(Y_{S}/X_{S})\neq P(Y_{T}/X_{T})$. For example, an assessment may be done on the topic of the resort in the $\mathbb{D}_{S}$ and it will be used to train a model for restaurants in the $\mathbb{D}_{T}$. DA is mostly effective when the $\mathbb{T}_{T}$ has a distinct distribution or there is a scarcity of labeled data.
(b) Cross-modality DTL: Also known as cross-lingual DTL in the spoken language field, most DTL methods, more or less, a connection in feature spaces or label spaces is required between $\mathbb{D}_{S}$ and $\mathbb{D}_{T}$. In other words, DTL can only occur when the source and destination data are both in the same modality, like video, speech, or text. cross-lingual DTL, in contrast to all other DTL approaches, is one of the most complicated issues in DTL. It is assumed that the feature spaces of the source and destination domains are completely distinct ($\chi_{S}\neq\chi_{T}$), as in speech-to-image, image-to-text, and text-to-speech. Furthermore, the label spaces of source $Y_{S}$ and destination $Y_{S}$ domains might differ ($Y_{S}\neq Y_{T}$).
(c) Unsupervised DTL: Intends to enhance the learning of the target predictive function $\mathbb{F}_{T}$ in $\mathbb{D}_{T}$ using the knowledge in $\mathbb{D}_{S}$ and $\mathbb{T}_{S}$, where $\mathbb{T}_{S}$ different from $\mathbb{T}_{T}$ but related, and $Y_{S}$ and $Y_{T}$ are not visible, given a SD $\mathbb{D}_{S}$ with a learning task $\mathbb{T}_{S}$, a TD $\mathbb{D}_{T}$ and a matching learning task $\mathbb{T}_{T}$ ($\mathbb{D}_{S}$ different from $\mathbb{D}_{T}$, but related).
2.1.2 Adversarial DTL
In contrast to the methods described above for DTL, adversarial learning [wang2020transfer] aids in the learning of more transferable and discriminative representations. The work in [ganin2016domain], was the first that introduced the domain-adversarial neural network (DANN). Instead of using a predefined distance function like maximum mean discrepancy (MMD), the core idea is to use a domain-adversarial loss in the network. This has greatly aided the network’s ability to learn more discriminative data. Many studies have used domain-adversarial training as a result of DANN’s idea [bousmalis2016domain, chen2019joint, long2017deep, zhang2018collaborative]. All of the previous work ignores the different effects of marginal and conditional distributions in adversarial TL, whereas in [wang2020transfer], the proposed scheme, named dynamic distribution alignment (DDA), can dynamically evaluate the importance of each distribution.

### v2

@@ Abstract
Recent advancements in deep learning (DL) have posed a significant challenge for automatic speech recognition (ASR). ASR relies on extensive training datasets, including confidential ones, and demands substantial computational and storage resources. Enabling adaptive systems improves ASR performance in dynamic environments. DL techniques assume training and testing data originate from the same domain, which is not always true. Advanced DL techniques like deep transfer learning (DTL), federated learning (FL), and deep reinforcement learning (DRL) address these issues. DTL allows high-performance models using small yet related datasets, FL enables training on confidential data without dataset possession, and DRL optimizes decision-making in dynamic environments, reducing computation costs.
This survey offers a comprehensive review of DTL, FL, and DRL-based ASR frameworks, aiming to provide insights into the latest developments and aid researchers and professionals in understanding the current challenges. Additionally, Transformers, which are advanced DL techniques heavily used in proposed ASR frameworks, are considered in this survey for their ability to capture extensive dependencies in the input ASR sequence. The paper starts by presenting the background of DTL, FL, DRL, and Transformers and then adopts a well-designed taxonomy to outline the state-of-the-art (SOTA) approaches. Subsequently, a critical analysis is conducted to identify the strengths and weaknesses of each framework. Additionally, a comparative study is presented to highlight the existing challenges, paving the way for future research opportunities.
@@ 1.1 Preliminary
Advancements in artificial intelligence (AI) have significantly improved human-machine interaction (HMI), especially with technologies that convert speech into executable actions. automatic speech recognition (ASR) emerges as a leading communication technology in HMI, extensively utilized by corporations and service providers for facilitating interactions through AI platforms like chatbots and digital assistants. Spoken language forms the core of these interactions, emphasizing the necessity for sophisticated speech processing in AI systems tailored for ASR. ASR technology encompasses the analysis of (i) acoustic, lexical, and syntactic aspects; and (ii) semantic understanding. The acoustic model (AM) processing includes speech coding [1], enhancement [2], source separation [3, 2], alongside securing speech via steganography [4, 5, 6] and watermarking [7, 8, 9]. These components are integral to audio analysis. On the other hand, the semantic model (SM), often identified as language model (LM) processing in literature, involves all natural language processing (NLP) techniques. This AI branch aims at teaching computers to understand and interpret human language, serving as the basis for applications like music information retrieval [10], sound file organization [11], audio tagging (AT), and event detection (ED) [12], as well as converting speech to text and vice versa [13], detecting hate speech [14], and cyberbullying [15]. Employing NLP across various domains enables AI models to effectively comprehend and respond to human inputs, unveiling extensive research prospects in diverse sectors.
@@
BERT
bidirectional encoder representations from transformers
@@
CLDNN
convolutional long-short term deep neural network
@@
DDPG
deep deterministic policy gradien
DDQN
double deep Q-network
@@
DSC
depthwise separable convolutions
DSLM
domain-specific language modeling
@@
GAN
generative adversarial network
GLDPT
global-local dual-path Transformer
HFL
horizontal federated learning
@@
HSD
hate speech detection
ISTFT
inverse short-time Fourier transform
LLM
large language model
@@
MMD
maximum mean discrepancy
MOS-LQO
mean opinion score-listening quality objective
MTL
multitask learning
@@
SARSA
state–action–reward–state–action
SCST
self-critical sequence training
@@
STFT
short-time Fourier tranform
SVD
singular value decomposition
@@
TSTNN
Transformer-based neural network
VFL
vertical federated learning
@@
Recent advancements in ASR have been significantly propelled by the evolution of deep learning (DL) methodologies. An extensive range of DL models has been developed, demonstrating remarkable improvements and surpassing former SOTA achievements [16, 17]. Transformers, a notable innovation within these DL approaches, have become a cornerstone in advancing various NLP tasks, including ASR. Initially conceptualized for sequence-to-sequence (S2S) applications in NLP, their success is largely attributed to their adeptness at discerning long-range dependencies and complex patterns within sequential data. A hallmark of Transformer models is their utilization of an attention mechanism, which precisely focuses on specific portions of the input sequence during prediction tasks. This mechanism is particularly effective in ASR, facilitating the detailed modeling of contextual nuances and the interconnections among acoustic signals, essential for accurate transcription. Models such as the Transformer Transducer, Conformer, and end-to-end speech processing network (ESPNet), leveraging self-attention and parallel processing, have achieved leading performance in ASR tasks. Their robustness across diverse languages further underscores their capability to adapt to a wide range of linguistic features and acoustic variations, making Transformers an exceptionally promising option for enhancing ASR systems, surpassing the constraints of conventional models.
@@ 1.2 Contribution of the paper
In recent years, numerous survey papers have been published to assess various aspects of ASR models. Some of these surveys concentrate on specific languages, such as Portuguese [23], Indian [24], Turkish [25], Arabic [26] and tonal languages (including Asian, Indo-European and African) [27]. Additionally, Abushariah et al.’s review emphasizes bilingual ASR [28]. On the non-specific language review front, specific areas within ASR have been targeted, including ASR using limited vocabulary [29], ASR for children [30], error detection and correction [31], and unsupervised ASR [32]. Systematic reviews with a focus on neural networks [33] and deep neural networks [34] have also been proposed. In another comprehensive review, Malik et al. [35] discussed diverse feature extraction methods, SOTA classification models, and some deep learning approaches. Recently, the authors presented an ASR review focused on DTL for ASR [36]. Table 1 presents a summary of the main contributions of the proposed DTL review compared to other existing DTL reviews/surveys.
This survey article offers several significant enhancements and additions compared to previous DTL surveys. Firstly, it consolidates works that utilize both DTL and advanced DL approaches, providing a comprehensive overview of their intersection. Secondly, it provides performance evaluation results of all considered approaches. Thirdly, it includes metrics and dataset reviews used in DTL models. Furthermore, it tackles ongoing challenges and consequently proposes future directions. The main contributions of this article can be summarized as follows:
@@
Table 1: Contribution comparison of the proposed contribution against other hand DTL review. The tick mark (✓) indicates that the specific field has been addressed, whereas the cross mark () means addressing the specific fields has been missed.
Refs Year Description of the survey/review Advanced DL methods Performances Metrics Dataset Current Future
DRL FL DTL Transf. evaluation review challe/Gaps directions
[31] 2018 ASR review for errors detection and correction
[34] 2019 Systematic review on DL-based speech recognition
[24] 2020 ASR survey for Indian languages
[23] 2020 ASR survey for Portuguese language
[25] 2020 ASR survey for Turkish language
[35] 2021 ASR survey
[27] 2021 ASR survey for tonal languages
[32] 2022 Unsupervised ASR review
[30] 2022 ASR Systematic review for children
[29] 2022 ASR survey for limited vocabulary
[26] 2022 ASR Systematic review for Arabic language
[28] 2022 Bilingual ASR review
[33] 2023 ASR survey on neural network techniques
[36] 2023 ASR based on DTL review
Our 2024 ASR review on advanced DL techniques
@@ 1.4 Structure of the paper
This paper is structured into six sections. The current section provides an introduction to the paper. Section 2 providing background on AM and LM, and reviewing evaluation metrics and datasets utilized in ASR. Moving forward, Section 3 delves into a comprehensive review of recent advancements in ASR utilizing advanced DL approaches, including Transformers, DTL, FL and DRL. Sections 4 and 5 respectively address the existing challenges and future directions concerning advanced DL-based ASR. Finally, Section 6 presents concluding remarks summarizing the key findings of the paper. Figure 2 presents a structured roadmap, offering a comprehensive guide to assist readers in navigating through the various sections and subsections of the paper.